# Kiet A. NguyenAdheesh JuvekarTianjiao YuMuntasir WahedIsmini Lourentzou

### v1

@@ Abstract
Recent advances in Large Vision-Language Models (LVLMs) have sparked significant progress in general-purpose vision tasks through visual instruction tuning. While some works have demonstrated the capability of LVLMs to generate segmentation masks that align phrases with natural language descriptions in a single image, they struggle with segmentation-grounded comparisons across multiple images, particularly at finer granularities such as object parts. In this paper, we introduce the new task of part-focused semantic co-segmentation, which seeks to identify and segment common and unique objects and parts across images. To address this task, we present Calico, the first LVLM that can segment and reason over multiple masks across images, enabling object comparison based on their constituent parts. Calico features two proposed components, a novel Correspondence Extraction Module, which captures semantic-rich information to identify part-level correspondences between objects, and a Correspondence Adaptation Module, which embeds this information into the LVLM to facilitate multi-image understanding in a parameter-efficient manner. To support training and evaluation, we curate MixedParts, a comprehensive multi-image segmentation dataset containing $\sim$2.4M samples across $\sim$44K images with diverse object and part categories. Experimental results show Calico, finetuned on only 0.3% of its architecture, achieves robust performance in part-focused semantic co-segmentation.
Figure 1: Multi-Image Part-focused Object Comparison with Calico. Our pixel-grounded Large Vision-Language Model, Calico, can perform part-focused semantic co-segmentation, a newly introduced task where the goal is to identify, segment, and label common objects, as well as common and unique object parts across multiple images.
@@ 1 Introduction
Analyzing objects by decomposing them into their constituent parts can enhance understanding of inter-object relationships both within and across categories. This part-level understanding is crucial for applications requiring detailed object comparisons, such as robotic manipulation, medical imaging, and educational tools. By recognizing similarities and differences at the part level across multiple images, these applications can enable more context-driven analysis and actions. Detailed comparisons that identify shared and unique parts offer insights into the critical features and functions of objects. For example, while both spoons and forks have handles, this part is not central to their primary functions. Instead, distinguishing the fork’s tines from the spoon’s bowl allows tasks like robotic grasping or visual comparisons to differentiate and interact with these objects based on their unique functions.
@@
Designing effective methods to analyze multiple images featuring diverse objects, and locate and identify their shared and distinct parts, presents an intriguing challenge. Given two images of similar objects, the goal is to generate segmentation masks for the shared or distinct parts (locate), establish one-to-one correspondences across images for comparative analysis (compare), and assign descriptive labels to these parts (identify). We term this task part-focused semantic co-segmentation.
@@
Research in part-focused semantic segmentation has explored various facets of this related granular task. Many studies have centered on single-image localized part learning [4, 58, 29, 38, 5, 63], focusing on segmenting parts within an object as a terminal task [4, 38] or as a means to support broader tasks like object or human parsing [53, 75]. However, these methods are not designed for multi-image comparisons, as the segmented parts from different images lack a consistent mapping of corresponding semantic components across images. Conversely, some research in this realm has tackled the related task of part co-segmentation, where models learn to segment objects across multiple images in a semantically consistent manner, despite variations in pose, shape, camera angle, etc. [20, 1, 33, 13, 5]. Yet, these approaches often require the number of part classes to be specified as input and struggle to identify unique parts between objects. Most part co-segmentation methods segment the entire object into parts, covering all areas of the object, which limits the model’s flexibility in pinpointing and analyzing subtle, part-specific variations that are essential for comparing different objects. Furthermore, the unsupervised or semi-supervised nature of recent methods hampers their ability to accurately label object parts.
@@
Recently, vision research has experienced a surge of interest in leveraging LVLMs for object segmentation tasks [24, 46, 50, 68, 60, 67]. Following LISA [24], many studies have adopted approaches that output additional segmentation tokens [46, 50, 68], which contain mask embeddings of objects within an image. These tokens are then processed by a decoder to generate the final segmentation masks. Such methods have proven highly effective for unifying tasks like semantic segmentation and referring expression segmentation within a single architecture, thanks to the adaptability of LVLMs in handling diverse input and output types. Although some studies have demonstrated varying levels of part understanding [24, 46, 67], they do not address reasoning about parts across multiple images or segmenting shared and distinctive parts between objects.
@@
To this end, we introduce Component-Focused Adaptive Learning for Multi-Image Co-Localization of Objects (Calico), a Vision-Language Model (VLM) designed to perform localized object comparison across image pairs. CALICO’s architecture is unique in its design, combining a Correspondence Extraction Module to capture part-level semantic relationships and a Correspondence Adaptation Module to embed these relationships efficiently into an LVLM. These components enable multi-image, part-level understanding without substantially increasing the parameter count. Calico augments a pretrained segmentation-based LVLM with a parameter-efficient adapter module to learn multi-image co-localization at multiple granularities. To extract semantic correspondence information between images, we propose a novel Correspondence Extraction Module, which employs a frozen encoder with strong semantic correspondence capabilities learned through self-supervised training [3]. This extracted information is then incorporated within the model using a Correspondence Adaptation Module applied at various layers of the pretrained LVLM to facilitate inter-image understanding. Due to the lightweight nature of these adaptive modules, the trainable parameters in Calico represent only 0.3% ($\sim$29M) of the entire architecture. To the best of our knowledge, Calico represents the pioneering effort in training an LVLM to perform multi-image part co-segmentation.
@@
To enable effective training for this novel task, we introduce the Multi-Image Cross-Segmentation of Distinctive and Common Parts (MixedParts) dataset that contains diverse and logically comparable object pairs, allowing Calico to generalize across varied categories and visual details. Leveraging widely used, publicly available part segmentation datasets [45, 73, 74, 61, 15], we manually curate pairs of object labels that are logically comparable and share at least one common part label. For instance, pairing a “chair” with an “ottoman” is logical since both belong to the category of seating furniture, making them more comparable than, e.g., a “chair” and a “microwave.” We then construct image pairs based on these labels to train the model to co-segment not only the objects themselves but also their shared and unique parts across multiple images.
@@
Lastly, as our work is the first to address multi-image part co-segmentation, we create baselines using publicly available pretrained models to tackle this novel task. We conduct extensive experiments on MixedParts to evaluate the effectiveness of our approach, including ablation studies to assess the contributions of the proposed correspondence extraction and adaptation modules. Calico achieves superior performance in part-focused semantic co-segmentation, showing a 20.8% relative mean IoU improvement on MixedParts compared to the next best baseline. In summary, our contributions are as follows:
@@
  • We introduce the novel task of part-focused semantic co-segmentation, which aims to co-segment and label common and unique parts between objects across multiple images for granular object comparison. To the best of our knowledge, this is the first work to formalize this multi-image object/part co-segmentation task.
@@
  • We propose Calico, an LVLM designed for part-focused semantic co-segmentation. Calico incorporates a novel correspondence extraction module and an adaptation module to learn semantic correspondences and localized co-segmentation across multiple images in a parameter-efficient manner.
@@
  • We introduce the MixedParts dataset for part-focused semantic co-segmentation, compiled from diverse part segmentation datasets and featuring images of logically comparable objects and parts.
@@
  • We construct baselines using publicly available pretrained models and conduct experiments to evaluate Calico’s performance on MixedParts, including ablation studies to analyze the contributions of the correspondence extraction and adaptation modules.
@@ 2 Related Work
2.1 Image Segmentation
Image segmentation, a fundamental task in computer vision, has garnered extensive research over the years, with seminal works [35, 16, 22] paving the way for modern approaches. Recently, SAM [23] has emerged as a leading segmentation method due to its zero-shot capabilities and ability to generate accurate masks for novel objects. Despite its strong zero-shot performance, SAM cannot identify masked objects given arbitrary text input and relies on region inputs to operate effectively. To address this limitation, Semantic-SAM [26] employs text encoder outputs atop its decoder to label segmentation masks, while also enabling labeled segmentation at any granularity. SEEM [77] introduces a joint image-text representation space allowing labeling of segmentation masks. SAM demonstrates robust segmentation capabilities even when combined with LVLMs [60, 46, 24]. Thus, we incorporate a SAM decoder in Calico. Building upon these works, we further fine-tune SAM’s pixel decoder using the LLM’s special token output embeddings to produce the desired segmentation masks. While the aforementioned works focus on segmenting individual images, Calico tackles multi-image object and part co-segmentation.
@@ 2.2 Part Segmentation
Part segmentation enables a more granular understanding by parsing objects into meaningful components or parts. While numerous approaches have been proposed to address this task, many are tailored to specific domains [66, 65, 21, 76] or are constrained to closed-set vocabularies [9, 29, 39, 38]. A recent approach, VLPart [56], efficiently tackles this task with open vocabulary predictions by utilizing a pretrained LVLM [44] and pretrained DINO features [3, 41] to perform semantic correspondence with base objects, enabling the model to label any part of the object. However, in practice, the labels are constrained by the input to the text encoder. Many recent works [24, 46, 50, 68, 60, 67] integrate the reasoning capabilities of LVLMs with segmentation models to augment visual understanding of the LVLMs. These models enable open-vocabulary object segmentation, without being constrained to the input prompt. Most of these models, motivated by LISA [24], utilize a pretrained LVLM such as LLaVA [32] in tandem with a decoder to predict segmentation masks.
While several of these works can also segment object parts [24, 46, 67] since their training data consists of large annotated datasets containing part information, their part segmentation capabilities often lag behind their object segmentation capabilities. Additionally, they also require explicit mention of object parts in the input prompt for successful segmentation. In other words, simply requesting to “segment multiple parts of the given object” does not yield the desired results. Instead, each part must be individually specified, e.g. by requesting to segment the leg of the chair to obtain the desired segmentation maps. In contrast, Calico augments a pretrained LVLM with object and part co-segmentation capabilities without the need for explicit part-specific prompts, simplifying user input requirements and allowing flexibility in instructing the model to infer and delineate various objects and parts across different images.
@@ 2.3 Object/Part Co-Segmentation
Figure 2: Overview of the Calico Architecture for Part-Focused Semantic Co-Segmentation. Calico uses a Q-Former cross-attention module to query efficient image embeddings from a pretrained image encoder, which are passed into a Vicuna-based LLM as image features. We extract [SEG] tokens from the output text, which are used to prompt a SAM decoder to output corresponding segmentation masks. We propose two modules, the Correspondence Extraction Module (CEM) and the Correspondence Adaptation Module (CAM), to learn semantic-rich features for multi-image correspondence information; details in Figure 3.
Co-segmentation aims to discover common objects across multiple images. Early co-segmentation works [25, 72, 69] employ CNNs with fully supervised training or fine-tuning on co-segmentation datasets. LCCo [11] leverages the semantic understanding of CLIP [44] to perform object co-segmentation. However, these approaches do not extend to co-segmenting object parts. Conversely, part co-segmentation involves simultaneously segmenting corresponding parts across multiple images. Previous methods [20, 1, 33, 13, 5] tackle this task with unsupervised or self-supervised learning. DFF [6] employs matrix factorization with features from a pretrained CNN, while SCOPS [20] trains an encoder-decoder CNN with an equivariance loss for robust part segmentation and a semantic consistency loss for improved co-segmentation. Moreover, a recent work [1] identifies that DINO features encapsulate semantic and spatial correspondences that can be utilized to find similar parts of objects across different images.
However, although these methods can co-segment parts, they lack the ability to predict semantic part labels and generate segmentation masks for unique object parts. In this work, we introduce Calico, a model that harnesses the zero-shot capabilities of SAM and the rich semantic information obtained from DINOv2 features [41, 8] for part co-segmentation in conjunction with part labeling. Calico integrates LVLMs with co-segmentation to improve object/ part reasoning and comparison across different images with diverse scenes, by predicting co-segmentation masks and semantic labels for both unique and common parts.
@@ 3.2 Calico Architecture
Calico is an LVLM designed to output multiple segmentation masks per image for a series of images that address commonalities and differences across images. In addition to its core functionality, our model incorporates modules aimed at integrating semantic correspondences between similar objects across images, alongside multi-image understanding and segmentation. As illustrated in Fig. 2, the architecture is composed of a Vicuna-based large language model $\mathcal{M}$ in tandem with a vision module $\mathcal{I}$ and a vision-to-language projection layer, which projects image embeddings from $\mathcal{I}$ into $\mathcal{M}$’s language space.
@@
The final input $\mathbf{T}^{0}\!\in\!\mathbb{R}^{S\times D}$ into $\mathcal{M}$ is composed of interleaved text and image tokens $,\mathbf{T}^{0}\!=\!\{\mathbf{t}^{0}_{1},\cdots,\mathbf{v}^{0}_{11},$$\cdots,\mathbf{v}^{0}_{1S_{I}},\cdots,$ $\mathbf{t}^{0}_{i},\cdots,\mathbf{v}^{0}_{j1},\cdots,\mathbf{v}^{0}_{jS_{I}},\cdots,\mathbf{t}^{0}_{S_{T}}\}$, where $\mathbf{t}^{0}_{i}$ is the $i^{\textnormal{th}}$ embedded text token at layer 00 ($1\leq i\leq S_{T}$), $\mathbf{I}^{0}_{j}\!=\!\left\{\mathbf{v}^{0}_{j1},\cdots,\mathbf{v}^{0}_{jS_{I}}\right\}$ are the $S_{I}$ tokens pertaining to the $j^{\textnormal{th}}$ image, and the superscript $k$ of $\mathbf{T}^{k}$ represents the LLM output at layer $k$ or input at layer $k+1$. The sequence length $S$ of $\mathbf{T}^{0}$ thus sums up the text and image lengths, i.e., $S\!=\!S_{T}+N_{I}\times S_{I}$. Finally, we obtain predicted outputs from the $N$-layered LLM by $\mathbf{\hat{T}}^{N}\!=\!\mathcal{M}\left(\mathbf{T}^{0}\right)$. In practice, we encourage the LLM to learn comprehensive multimodal understanding through prompts such as “The <image> (IMAGE1) and <image> (IMAGE2) provide an overview of the pictures. Can you segment the common object in these images?”, where the <image> tokens are replaced with the projected embeddings of the corresponding image. Each image is also associated with a unique identifier (e.g., IMAGE1, IMAGE2) for more convenient and clear reference, avoiding any potential ambiguity from using ordinal terms to refer to individual images in multi-image settings, e.g., “the first” or “the second” image.
@@
Pixel-Grounded Outputs. To enable pixel-level grounding, we augment the model’s vocabulary with the segmentation token [SEG] and teach it to output grounding tags <p> and </p>, following recent work [46]. Through supervision, the model learns to ground a noun phrase associated with the following segmentation token by enclosing it in the grounding tags, which immediately precede the corresponding [SEG] token (e.g., The unique parts of the objects are <p> the seat cushion </p> [SEG] (IMAGE1) and <p> the back pillow </p> [SEG] (IMAGE2).”). We append the image identifiers immediately following [SEG] tokens to distinguish between tokens belonging to different images.
To transform the [SEG] tokens $\mathbf{S}\!=\!\left\{\mathbf{S}_{1},\!\cdots,\!\mathbf{S}_{N_{I}}\right\}\subset\mathbf{T}^{K}$, where each $\mathbf{S}_{i}\in\mathbb{R}^{S_{j}\times D}$ corresponds to the set of $S_{j}$ predicted masks associated with image $i$, into segmentation mask sets $\mathbf{\hat{M}}\!=\!\left\{\mathbf{\hat{M}}_{1},\cdots,\mathbf{\hat{M}}_{N_{I}}\right\}$, our architecture incorporates a Transformer-based grounding model [23], composed of a grounding encoder $\mathcal{G}$ and a pixel decoder $\mathcal{D}$. The input images $\mathbf{X}_{\textnormal{image}}$ are passed through the frozen encoder to obtain the grounding embeddings $\mathbf{X}_{\textnormal{ground}}$ as the vision signal for the decoder by $\mathbf{X}_{\textnormal{ground}}\!=\!\mathcal{G}\left(\mathbf{X}_{\textnormal{image}}\right)\!\in\!\mathbb{R}^{N_{I}\times S_{D}\times D_{D}}$. For an encoded image $\mathbf{X}_{\text{ground }i}$, the tokens in $\mathbf{S}_{i}$ act as prompts for the finetuned pixel decoder after being passed through a projection layer $f_{\text{segmentation}}:\mathbb{R}^{D}\mapsto\mathbb{R}^{D_{D}}$. Finally, the decoder $\mathcal{D}$ produces binary segmentation masks accordingly by:

### v2

@@ Abstract
Recent advances in Large Vision-Language Models (LVLMs) have enabled general-purpose vision tasks through visual instruction tuning. While existing LVLMs can generate segmentation masks from text prompts for single images, they struggle with segmentation-grounded reasoning across images, especially at finer granularities such as object parts. In this paper, we introduce the new task of part-focused semantic co-segmentation, which involves identifying and segmenting common objects, as well as common and unique object parts across images. To address this task, we present Calico, the first LVLM designed for multi-image part-level reasoning segmentation. Calico features two key components, a novel Correspondence Extraction Module that identifies semantic part-level correspondences, and Correspondence Adaptation Modules that embed this information into the LVLM to facilitate multi-image understanding in a parameter-efficient manner. To support training and evaluation, we curate MixedParts, a large-scale multi-image segmentation dataset containing $\sim$2.4M samples across $\sim$44K images spanning diverse object and part categories. Experimental results demonstrate that Calico, with just 0.3% of its parameters finetuned, achieves strong performance on this challenging task.
Figure 1: Multi-Image Part-focused Object Comparison with Calico. Our pixel-grounded Large Vision-Language Model, Calico, performs part-focused semantic co-segmentation, a newly introduced task where the goal is to identify, segment, and label common objects, as well as common and unique object parts across multiple images.
@@ 1 Introduction
Analyzing objects by decomposing them into their constituent parts can enhance understanding of inter-object relationships both within and across categories. This part-level reasoning is crucial for applications requiring detailed object comparisons, such as robotic manipulation, medical imaging, and educational tools. Detailed comparisons that identify shared and unique parts offer insights into the critical features and functions of objects. For example, while both spoons and forks have handles, this part is not central to their primary functions. Instead, distinguishing the fork’s tines from the spoon’s bowl allows tasks like robotic grasping or visual comparisons to differentiate and interact with these objects based on their unique functions.
@@
Designing effective methods to analyze multiple images featuring diverse objects by locating and identifying their shared and distinct parts presents an intriguing challenge. Given a pair of images containing similar objects, the goal is to generate segmentation masks for the shared or distinct parts (locate), establish one-to-one correspondences across images for comparative analysis (compare), and assign descriptive labels to the discovered parts (identify). We term this task part-focused semantic co-segmentation.
@@
Existing research has explored various facets of this related granular task. Many works have investigated localized part learning [6, 62, 33, 42, 7, 67], focusing on segmenting parts within an object as a terminal task [6, 42] or as a means to support broader tasks like object or human parsing [57, 79]. However, these methods are not designed for multi-image comparisons, as the segmented parts from different images lack a consistent mapping of corresponding semantic components across images. Conversely, some research in this realm has tackled the related task of part co-segmentation, where models learn to segment objects across multiple images in a semantically consistent manner, despite variations in pose, shape, camera angle, etc. [23, 2, 37, 16, 7]. Yet, these approaches often require the number of part classes to be specified as input and struggle to identify unique parts between objects.
@@
Recent work has shown a growing interest in leveraging LVLMs for object segmentation tasks [27, 50, 54, 72, 64, 71]. Following LISA [27], many methods introduce additional segmentation tokens [50, 54, 72], which contain mask embeddings of objects within an image. These tokens are then processed by a decoder to generate the final segmentation masks. Such methods have proven highly effective for unifying tasks like semantic segmentation and referring expression segmentation within a single architecture, thanks to the adaptability of LVLMs in handling diverse input and output types. While some of these models exhibit partial understanding of object parts [27, 50, 71], they are not designed for reasoning over parts across multiple images or for segmenting shared and unique parts between objects.
@@
To this end, we introduce Component-Focused Adaptive Learning for Multi-Image Co-Localization of Objects (Calico), a vision-language model designed to perform localized object comparison across image pairs. To the best of our knowledge, Calico is the first LVLM trained for multi-image part-level co-segmentation. Calico augments a pretrained segmentation-based LVLM with parameter-efficient adapter modules to learn multi-image co-localization at multiple granularities. To extract part-level semantic correspondence information between images, we propose a novel Correspondence Extraction Module, which employs a frozen DINOv2 encoder with strong semantic correspondence capabilities learned through self-supervised training [5]. These correspondences are then injected into the pretrained LVLM via Correspondence Adaptation Modules applied at selected layers to facilitate inter-image understanding. Due to the lightweight nature of these adaptive modules, the trainable parameters in Calico represent only 0.3% ($\sim$29M) of the entire architecture.
@@
To support training for this novel task, we introduce the Multi-Image Cross-Segmentation of Distinctive and Common Parts (MixedParts) dataset that contains diverse and logically comparable object pairs that enable Calico to generalize across varied categories and visual details. Building on widely used, publicly available part segmentation datasets [49, 77, 78, 65, 18], we manually curate pairs of object classes that are logically comparable and share at least one common part. For instance, pairing a “chair” with an “ottoman” is meaningful since both belong to the category of seating furniture, making them more comparable than, e.g., a “chair” and a “microwave.” We then assemble image pairs based on these class pairings, enabling training for co-segmentation of not only the objects themselves but also their shared and unique parts across multiple images.
Figure 2: Calico Efficiency. Calico improves performance (e.g., recall) in part-focused co-segmentation while reducing TFLOPS by $\sim$32-35and accelerating inference by $\sim$30-51compared to SotA baselines, using 8-18$\times$ fewer image tokens.
@@
Since our work is the first to address part-focused semantic co-segmentation, we construct baselines by adapting publicly available pretrained models to this novel task. We conduct extensive experiments on MixedParts to evaluate the effectiveness of Calico, including ablations to quantify the contributions of the proposed correspondence extraction and adaptation modules. Calico outperforms all baselines on part-focused semantic co-segmentation, achieving a 6.3% relative improvement in mean IoU on MixedParts over the next best method, while using 18 times fewer image tokens, reducing TFLOPS by 32.6%, and accelerating inference by 51.3% (as shown in Figure 2). In summary, the contributions of our work are as follows:
[…]
@@
    We introduce the novel task of part-focused semantic co-segmentation, which involves co-segmenting and labeling common and unique parts between objects across images for granular object comparison. To the best of our knowledge, this is the first work to formalize this multi-image object/part co-segmentation task.
[…]
@@
    We propose Calico, an LVLM for part-focused semantic co-segmentation. Calico incorporates a novel Correspondence Extraction Module and parameter-efficient Adaptation Modules to learn semantic correspondences and enable efficient co-segmentation.
[…]
@@
    We curate MixedParts, a large-scale benchmark for part-focused semantic co-segmentation, featuring logically comparable objects and parts drawn from publicly available diverse part segmentation datasets.
[…]
@@
    We establish strong baselines from publicly available pretrained models and conduct experiments to evaluate Calico’s performance on MixedParts, including comprehensive ablations to analyze the contributions of our proposed modules.
@@ 2 Related Work
2.1 Prompting Image Segmentation
Image segmentation, a fundamental computer vision task, has garnered extensive research over the years, with seminal works [39, 19, 25] paving the way for modern approaches. Pretrained on large-scale data, SAM [26] has emerged as a leading foundation model for image segmentation, capable of interpreting diverse prompts and generating corresponding masks. However, despite its strong zero-shot performance, SAM lacks semantic labeling, limiting its practical use. SEEM [81] addresses this by introducing a joint image-text representation space inspired by LVLMs, enabling unified prompting and effective mask labeling. Still, both SAM and SEEM struggle with fine-grained understanding, such as object-part segmentation. Semantic-SAM [29] extends SAM by incorporating multi-level image embeddings tied to different prompt types for multi-granularity understanding but, like SAM, lacks text labels. Recent works [64, 50, 27, 54] combine LVLMs with SAM to bridge language and mask understanding, achieving more robust segmentation and labeling capabilities. Building on these advances, Calico integrates SAM with a finetuned pixel decoder to output segmentation masks with text labels. Unlike prior work focused on single-image segmentation, Calico also enables multi-image multi-granularity understanding through object and part co-segmentation.
@@ 2.2 Part Segmentation
Part segmentation enables fine-grained understanding by decomposing objects into semantically meaningful parts. While numerous approaches have been proposed to address this task, many are tailored to specific domains [70, 69, 24, 80] or are constrained to closed-set vocabularies [12, 33, 43, 42]. A recent approach, VLPart [60], tackles open vocabulary part segmentation by utilizing a Mask R-CNN backbone [19] trained via contrastive learning to align predicted masks with CLIP text features [48], enabling the model to label any object-part. PartGLEE [32] adopts a Q-Former-based architecture [30] to achieve similar alignment of predicted masks with CLIP text features; however, in practice, the labels are constrained by the input to these models’ text encoders. Many recent works [27, 50, 54, 72, 64, 71] integrate the reasoning capabilities of LVLMs with segmentation models to augment visual understanding of the LVLMs. These models enable open-vocabulary object segmentation without being constrained to the input prompt. Most of these models, motivated by LISA [27], utilize LLaVA-based LVLMs  [36] in tandem with a decoder to predict segmentation masks.
While several of these works can also segment object parts [27, 50, 71] – since their training data consists of large annotated datasets with part information – their part segmentation capabilities often lag behind their object segmentation capabilities. Additionally, they also require explicit mention of object parts in the input prompt for successful segmentation and fail to segment multiple parts of an object. Instead, each part must be individually specified, e.g., by requesting to “segment the leg of the chair,” to obtain the desired segmentation maps. In contrast, Calico augments a pretrained LVLM with object and part co-segmentation capabilities without the need for explicit part-specific prompts, simplifying user input requirements and allowing flexibility in instructing the model to infer and delineate various objects and parts across different images.
@@ 2.3 Object/Part Co-Segmentation
Figure 3: Overview of the Calico Architecture for Part-Focused Semantic Co-Segmentation. Calico employs a Q-Former cross-attention module to query efficient image embeddings from a pretrained image encoder, which are passed as visual tokens into a Vicuna-based LLM. We extract [SEG] tokens from the output text, which are used to prompt a SAM decoder to produce corresponding segmentation masks. We propose two modules: the Correspondence Extraction Module (CEM), which captures semantic-rich part correspondences, and Correspondence Adaptation Modules (CAMs), which inject this information into the LVLM. CEM/CAM details in Figure 4.
Co-segmentation aims to identify and segment common objects across multiple images. Early works [28, 76, 73] employ CNNs with fully supervised training or finetuning on co-segmentation datasets, while later works leverage the semantic understanding of CLIP [14] or finetune an LVLM on samples containing similar objects [47]. However, these methods do not extend to co-segmenting object parts. Conversely, part co-segmentation involves simultaneously segmenting corresponding object parts across multiple images. Previous methods tackle this task with unsupervised or self-supervised learning [23, 2, 37, 16, 7]. DFF [8] employs matrix factorization on features from a pretrained CNN, while SCOPS [23] trains an encoder-decoder CNN with equivariance and semantic consistency losses for improved part co-segmentation. More recently, [2] shows that DINO features capture semantic and spatial correspondences useful for identifying similar parts across images.
While effective at discovering part-level correspondences, these methods lack the ability to assign semantic part labels or generate segmentation masks for unique parts. In this work, we introduce Calico, a model that harnesses SAM’s segmentation capabilities and DINOv2’s semantic features [45, 11] to jointly perform part co-segmentation and labeling. By integrating LVLMs with co-segmentation, Calico improves object-part reasoning across images with diverse scenes, generating both co-segmentation masks and semantic labels for unique and common parts.
@@ 3.2 Calico Architecture
Calico is an LVLM designed to generate multiple segmentation masks per image across a series of images, highlighting both shared and distinct regions among them. In addition to its core functionality, our model incorporates modules aimed at integrating semantic correspondences between similar objects across images, alongside multi-image understanding and segmentation. As illustrated in Figure 3, the architecture is composed of a Vicuna-based LLM $\mathcal{M}$ in tandem with a vision module $\mathcal{I}$ and a vision-to-language projection layer, which projects image embeddings from $\mathcal{I}$ into $\mathcal{M}$’s language space.
@@
The final input $\mathbf{T}^{0}\!\in\!\mathbb{R}^{S\times D}$ into $\mathcal{M}$ is composed of interleaved text and image tokens $,\mathbf{T}^{0}\!=\!\{\mathbf{t}^{0}_{1},\cdots,\mathbf{v}^{0}_{11},$$\cdots,\mathbf{v}^{0}_{1S_{I}},\cdots,$ $\mathbf{t}^{0}_{i},\cdots,\mathbf{v}^{0}_{j1},\cdots,\mathbf{v}^{0}_{jS_{I}},\cdots,\mathbf{t}^{0}_{S_{T}}\}$, where $\mathbf{t}^{0}_{i}$ is the $i^{\textnormal{th}}$ embedded text token at layer 00 ($1\leq i\leq S_{T}$), $\mathbf{I}^{0}_{j}\!=\!\left\{\mathbf{v}^{0}_{j1},\cdots,\mathbf{v}^{0}_{jS_{I}}\right\}$ are the $S_{I}$ tokens pertaining to the $j^{\textnormal{th}}$ image, and the superscript $k$ of $\mathbf{T}^{k}$ represents the LLM output at layer $k$ or input at layer $k+1$. The sequence length $S$ of $\mathbf{T}^{0}$ thus sums up the text and image lengths, i.e., $S\!=\!S_{T}+N_{I}\times S_{I}$. Finally, we obtain predicted outputs from the $N$-layered LLM by $\mathbf{\hat{T}}^{N}\!=\!\mathcal{M}\left(\mathbf{T}^{0}\right)$. In practice, we encourage the LLM to develop comprehensive multimodal understanding through prompts such as “The <image> (IMAGE1) and <image> (IMAGE2) provide an overview of the pictures. Can you segment the common object in these images?”, where the <image> tokens are replaced with the projected embeddings of the corresponding image. Each image is also associated with a unique identifier (e.g., IMAGE1, IMAGE2) for more convenient and clear reference, avoiding any potential ambiguity from using ordinal terms to refer to individual images in multi-image settings, e.g., “the first” or “the second” image.
@@
Pixel-Grounded Outputs. To enable pixel-level grounding, we augment the model’s vocabulary with the segmentation token [SEG] and teach it to output grounding tags <p> and </p>, following recent work [50]. Through supervision, the model learns to ground a noun phrase associated with the following segmentation token by enclosing it in the grounding tags, which immediately precede the corresponding [SEG] token (e.g., “The unique parts of the objects are <p> the seat cushion </p> [SEG] (IMAGE1) and <p> the back pillow </p> [SEG] (IMAGE2).”). We append the image identifiers immediately following [SEG] tokens to distinguish between tokens belonging to different images. To transform the [SEG] tokens $\mathbf{S}\!=\!\left\{\mathbf{S}_{1},\!\cdots,\!\mathbf{S}_{N_{I}}\right\}\subset\mathbf{T}^{K}$, where each $\mathbf{S}_{i}\in\mathbb{R}^{S_{j}\times D}$ corresponds to the set of $S_{j}$ predicted masks associated with image $i$, into segmentation mask sets $\mathbf{\hat{M}}\!=\!\left\{\mathbf{\hat{M}}_{1},\cdots,\mathbf{\hat{M}}_{N_{I}}\right\}$, our architecture incorporates a Transformer-based grounding model [26], composed of a grounding encoder $\mathcal{G}$ and a pixel decoder $\mathcal{D}$. The input images $\mathbf{X}_{\textnormal{image}}$ are passed through the frozen encoder to obtain grounding embeddings $\mathbf{X}_{\textnormal{ground}}$ as the vision signal for the decoder by $\mathbf{X}_{\textnormal{ground}}\!=\!\mathcal{G}\left(\mathbf{X}_{\textnormal{image}}\right)\!\in\!\mathbb{R}^{N_{I}\times S_{D}\times D_{D}}$. For an encoded image $\mathbf{X}_{\text{ground }i}$, the tokens in $\mathbf{S}_{i}$ act as prompts for the finetuned pixel decoder after being passed through a projection layer $f_{\text{segmentation}}\!:\!\mathbb{R}^{D}\mapsto\mathbb{R}^{D_{D}}$. Finally, the decoder $\mathcal{D}$ produces binary segmentation masks accordingly by: